OpenStudy (anonymous):

Solve for x: 2/3 (x-2) = 4x Options... -2/5 -5/2 2/5 5 Thank you:)

OpenStudy (ivettef365):

2/3(x-2) = 4x 2/3x - 4/3 = 4x 2/3x - 4x = 4/3

OpenStudy (ivettef365):

2x 4x -- - --- = 4/3 3 1

OpenStudy (ivettef365):

10x 4 - ---- = -- = -10x = 4 3 3

OpenStudy (ivettef365):

x = - 4 / 10 x = -2/5